# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

### v1

@@ 2.3 McCann’s results for 2-body systems
In this subsection, we review a slightly modified version of McCann’s results for Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizers and existence theorem of binary star solution (further discussion on modifications will be discussed in Section 3). We also make some remarks on them.
@@
    for some Lagrange multiplier $\lambda_{i}<0$ depending on the component;
@@
    the equations (EL) continue to hold on a $\delta$-neighbourhood of the $\Omega_{i}$;
@@
    if $P(\rho)$ satisfies assumption (F4’), then $\rho$ satisfies the reduced Euler-Poisson equations (EP’), where the pressure $P(\rho)$ is a function of the density under assumptions (F1)(F2)(F3).
@@
One might wonder whether $\rho$ in Theorem 2.25 (ii) is rotating. In fact, we can construct a truly rotating solution from it. Due to the discussion in subsection 2.2, if $\widetilde{\rho}$ is a local minimizer over ${R}\left(\mathbb{R}^{3}\right)$ of $E_{J}(\rho)$ — or by Corollary 4.9, even just over ${R}_{0}(\mathbb{R}^{3})$ — then we may define:
@@
where $\omega=\frac{J}{I\left(\rho\right)}$, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating binary star system (or star-planet system) such that $E(\rho,v)$ and $E_{J}(\rho)$ is local minimum. Moreover, $({\rho}(t,x),v(t,x))$ solves (EP) with $V(t,x)=V_{\widetilde{\rho}}\left(R_{-\omega t}x\right)$.
@@
In McCann’s paper [31], he uses the terminology “chemical potential” to call the “Lagrange multiplier” $\lambda_{i}$ in Parts (iv) and (v) here. Although it seems no chemical reaction occurs when stars move or rotate or are at rest, the origin of the term “chemical potential” can trace back to “the Lagrange multiplier $\mu=\frac{\partial E}{\partial N}$ in the Euler–Lagrange variational equation for the ground state density as the chemical potential of the electrons” [7]. We notice by Part (ii) that $\rho$ is a minimizer on some set with some restriction. Therefore, another way to understand this term is to identify “Lagrange multiplier” with the derivative of energy by (generalized) Lagrange multiplier theorem Theorem A.4 and Remark A.5, thus it can be called “potential”. Just as the derivative of gravitational potential energy is referred to as gravitational potential (up to sign).
@@
Notice in Theorem 2.25, since the local minimizer $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\bar{x}(\rho)=(0,0,0)^{T}$ in (EL). But when McCann tried to show the constrained minimizer $\rho^{J}$ (i.e. $J$ is fixed) is a Wasserstein $L^{\infty}$ local energy minimizer, he did not always require the center of mass of $\rho^{J}$ to be at original point. Therefore, we retain the term $\bar{x}(\rho)$ here to be consistent with the form of the Euler-Lagrange equations in McCann’s paper [31, Section 6].
@@
If we further assume the fluid particles move with bounded velocities, then $Y_{t}(y)$ will be a Lipshitz function of $t$ uniformly in $y$, and it is evident that (25) will be controlled by a multiple of |st||s-t|. Thus $\rho_{t}$ is a continuous function with respect to time $t$.
@@
On the other hand, if the topology on ${V}\left(\mathbb{R}^{3}\right)$ is such that the map taking $\boldsymbol{w}\in\mathbb{R}^{3}$ to $v(x):=\boldsymbol{w}\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous, then a converse to the statement above is also true: any local minimizer subject to the vector constraint $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$ also minimizes locally among the larger class of competitors with prescribed $J_{z}(\rho,v)=J$ [31, Remark 2.3]. This condition that $w\in\mathbb{R}^{3}\mapsto v(x):=w\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous may initially appear to have no obvious use. But we can consider the following arguments: due to [31, Remark 3.3], we can have an similar result to 2.17: if $\boldsymbol{J}$ is prescribed, the velocity which minimizes the energy is given by $\boldsymbol{v}(\rho,\boldsymbol{J})(x)=\boldsymbol{w}(\rho,\boldsymbol{J})\times x$, where $\boldsymbol{w}(\rho,\boldsymbol{J})\in\mathbb{R}$ is the unique angular velocity compatible with the given density $\rho$ and angular momentum $\boldsymbol{J}$. Notice the axis $\boldsymbol{w}$ of rotation may not coincide with the $z$-axis. One can show when $\rho$ is sufficiently close to $\rho^{\prime}$ in the sense of Wasserstein $L^{\infty}$ distance, $w(\rho,\boldsymbol{J})$ is sufficiently close to $w^{\prime}(\rho^{\prime},\boldsymbol{J^{\prime}})$. This condition that $w\in\mathbb{R}^{3}\mapsto v(x):=w\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous ensures that $\boldsymbol{v}(\rho,\boldsymbol{J})$ is sufficiently close to $\boldsymbol{v^{\prime}}(\rho^{\prime},\boldsymbol{J^{\prime}})$ in $V(\mathbb{R}^{3})$. In another word, local perturbations in ${R}_{0}\left(\mathbb{R}^{3}\right)$ can create local perturbations in ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$, which can help us to prove that any local minimizer subject to the vector constraint $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$ also minimizes locally among the larger class of competitors with prescribed $J_{z}(\rho,v)=J$, using similar arguments to [31, Remark 2.3 and Section 3].
@@
Now we can obtain a solution to (EP’) from the Euler-Lagrange equation (EL) thanks to Theorem 2.25 (iv). Conversely, for $\rho\in R_{0}(\mathbb{R}^{3})$ with N connected components to be a solution of (EP’), similar arguments as in subsection 2.2 shows that (EL) in Theorem 2.25 (iv) is necessary as we discuss in Remark 2.24.
By the arguments of Auchmuty and Beals [5] or Li [26] or McCann [31] we know a constrained energy minimizer on $W_{m,J}$ exists and its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. Then such constrained energy minimizer is indeed a local minimizer due to Lemma 4.7 (ii), and hence a solution to (EP’) after translation.
@@
What’s more, after another translation the center of mass of $\widetilde{\rho}$ can be 0 and is a local minimizer of $E_{J}(\rho)$. Let $v(x):=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\widetilde{\rho})}$, then the pair $(\widetilde{\rho},v)$ minimizes $E(\rho,v)$ locally on ${R}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ (thus on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right))$ subject to the constraint $J_{z}(\rho,v)=J$ or $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$. $\widetilde{\rho}$ satisfies reduced Euler-Poisson equations (EP’). Moreover, the uniformly rotating fluid $({\rho}(t,x),v(t,x))$ solves (EP) with $V(t,x)=V_{\widetilde{\rho}}\left(R_{-\omega t}x\right)$, here $({\rho}(t,x),v(t,x))=\left(\widetilde{\rho}\left(R_{-\omega t}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
@@
We point out that in McCann’s proof of Theorem 2.25, the discussion on how to transition from (EL) to (EP’) is not particularly detailed. One can see intuitively from subsection 2.2 that a refined discussion on differentiability is required. Therefore, in this section, we will go through the proof of Theorem 2.25, give new ideas to establishing differentiability (Part (vi) and Part (vii) of Theorem 2.25), and explain the differences between our proof and McCann’s.
@@
Given $s>0,h\in\mathbb{R}$ with $0<|h|\leq s$, let $t=\phi(s)$ and $t+l=\phi(s+h)$, then $h\rightarrow 0$ is equivalent to $l\rightarrow 0$ due to the continuity and monotonicity of $\phi$, and we also have $A^{\prime}(t)=s,A^{\prime}(t+l)=s+h$ since $\phi=\left(A^{\prime}\right)^{-1}$. Recall in Remark 2.5 we have $A^{\prime}(s)s-A(s)=P(s)$. Therefore, we have
@@
Let $B(l)=lA^{\prime}(t)-A(t+l)$, then $B$ is differentiable and $B^{\prime}(l)=A^{\prime}(t)-A^{\prime}(t+l)\neq$ 0 if $l\neq 0$, since $A^{\prime}$ is strictly increasing. Notice $lA^{\prime}(t)+A(t)-A(t+l)=B(l)-B(0)$, by Cauchy’s mean value theorem [35, 5.9 Theorem], $\exists\theta\in(0,1)$, such that $(B(l)-B(0))A^{\prime\prime}(t+\theta l)=\left(A^{\prime}(t+l)-A^{\prime}(t)\right)B^{\prime}(\theta l)$. Since $l\neq 0$, $\left(A^{\prime}(t+l)-A^{\prime}(t)\right)\neq 0$, $B^{\prime}(\theta l)\neq 0$, then both (B(l)B(0))(B(l)-B(0)) and $A^{\prime\prime}(t+\theta l)$ are also not 0. Then
@@
By Lagrange’s mean value theorem [35, 5.8 Theorem], $\exists\widetilde{\theta}\in(0,1)$, such that $A^{\prime}(t)-A^{\prime}(t+\theta l)=(-\theta l)A^{\prime\prime}(t+\widetilde{\theta}\theta l)$. Notice $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$ when $s>0$. Thus
@@
By Taylor’s theorem [35, 5.15 Theorem], we know $\exists\theta_{1}\in(0,1),\theta_{2}\in(0,1)$, $P^{\prime}(t+\theta l)=P^{\prime}(t)+\theta lP^{\prime\prime}(t)+\frac{(\theta l)^{2}}{2}P^{(3)}(t)+\cdots+\frac{(\theta l)^{n}}{n!}P^{(n+1)}\left(t+\theta_{1}\theta l\right)=\frac{(\theta l)^{n}}{n!}P^{(n+1)}$ $\left(t+\theta_{1}\theta l\right)\neq 0$ due to (F4’). Similarly, we have $P^{\prime}(t+\widetilde{\theta}\theta l)=\frac{(\widetilde{\theta}\theta l)^{n}}{n!}P^{(n+1)}\left(t+\theta_{2}\widetilde{\theta}\theta l\right)\neq 0$. Thus $\left|\frac{P^{\prime}(t+\widetilde{\theta}\theta l)}{P^{\prime}(t+\theta l)}\right|\leq 2$ when $|l|$ small enough. As $l\rightarrow 0$, we have $\frac{A^{\prime}(t)-A^{\prime}(t+\theta l)}{A^{\prime\prime}(t+\theta l)}\rightarrow 0$, and then $\frac{P\circ\phi(s+h)-P\circ\phi(s)}{h}\rightarrow t=\phi(s)$ as $h\rightarrow 0$. Thus, the derivative of $P\circ\phi$ is $\phi(s)$. By the same arguments we know $P\circ\phi$ is right differentiable at 0 with right derivatives $\phi(0)=$ 0. ∎
Remark 3.2.
Since (F4) implies (F4’), thus the result in Lemma 3.1 also holds true if $P$ satisfies (F1)(F2)(F3)(F4). One can also apply Cauchy’s mean value theorem directly to $\frac{P\circ\phi(s+h)-P\circ\phi(s)}{h}=\frac{P(t+l)-P(t)}{A^{\prime}(t+l)-A^{\prime}(t)}$ and obtain $\frac{P\circ\phi(s+h)-P\circ\phi(s)}{h}=\frac{P(t+l)-P(t)}{A^{\prime}(t+l)-A^{\prime}(t)}=\frac{P^{\prime}(t+\theta l)}{A^{\prime\prime}(t+\theta l)}=t+\theta l\rightarrow t=\phi(s)$ as $h\rightarrow 0$. We can do this because (F4), together with $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$ if $s>0$, implies $A^{\prime\prime}(t+\theta l)\neq 0$.
We skip the proof of Part (i) and Part (viii) of Theorem 2.25, one can refer to McCann’s paper [31]. Part (ii) and Part (iii) are closely related to the context, so we revisit McCann’s proof here. The other parts of the proof are relatively new, either addressing details not mentioned by McCann (Part (iv-v)) or providing new approachs to proving differentiability and explaining the necessity of modifying the original statements of the conclusions (Part (vi-vii)).
@@
Due to Part (ii), it suffices to consider $\rho$ that locally minimizes $E_{J}(\rho)$ on $R(\mathbb{R}^{3})$. We will give proofs of “weak” versions of Parts (iv-v) (Proposition 5.8, Proposition 5.11 and Remark 5.12) later in subsection 5.2, that is, (EL) holds for almost all points in the region we consider with negative $\lambda_{i}$. Based on this “weak” version of (v), we obtain the global continuity in Part (iii) as the following: first notice that almost everywhere in the region $\{\rho>0\}$, (EL) holds true and $A^{\prime}(\rho)>0$ implies $\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}>0$, thus on $\{\rho>0\}$, $A^{\prime}(\rho(x))=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}\leq V_{\rho}(x)+C$ since $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ has compact support. This result, together with Proposition A.7 (Hardy-Littlewood-Sobolev Inequality) or Proposition A.9, can help to show the bound of $\|\rho\|_{L^{p}}$ and $\left\|V_{\rho}\right\|_{L^{\widetilde{p}}}$ with $p$ and $\widetilde{p}$ increasing alternately and reaching infinity in finite steps. (Bootstrap method, similar as [5, Lemma 3 and Theorem A]). Then $V_{\rho}$ is continuously differentiable thanks to Proposition A.11. Then on a $\delta$-neighbourhood of $\{\rho>0\}$, denoted by $\Omega_{\delta}$, which contains $\partial\{\rho>0\}$, we have $\rho=\phi\circ\left[W_{\rho}\right]_{+}$ a.e., where $W_{\rho}(x)=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i},\phi=\left(A^{\prime}\right)^{-1}$ exists and continuous by Remark 2.7. It means $\rho$ has a continuous representative there. Notice the points in $\overline{\mathbb{R}^{3}\backslash\Omega_{\delta}}$ are the interior points of $\{\rho=0\}$, hence $\rho$ is also continuous in $\overline{\mathbb{R}^{3}\backslash\Omega_{\delta}}$. Thus, we conclude Part (iii).
@@
For Part (vi), notice in the region $\{\rho>0\}$, we have $\rho=\phi\circ W_{\rho}$ by Remark 2.26. Thanks to Part (iii) and $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\rho\in C_{c}^{0}\left(\mathbb{R}^{3}\right)$, where $C_{c}^{0}\left(\mathbb{R}^{3}\right)$ denotes the family of continuous functions with compact support. In particular, $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{q}\left(\mathbb{R}^{3}\right)$ for any $q>3$. By Proposition A.11, we know $V_{\rho}$ is continuous differentiable in $\mathbb{R}^{3}$. Since $P(\rho)$ satisfies assumption (F4), by Remark 2.8 we know $\phi\in C^{1}((0,\infty))$. Apply chain rule (one can refer to [38, Theorem 10.1.15]), we obtain $\rho\in C^{1}(\{\rho>0\})$. ∎

### v2

@@ 2.3 McCann’s results for 2-body systems
In this subsection, we revisit a slightly modified version of McCann’s results for Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizers and existence theorem of binary star solution (further discussion on modifications will be discussed in Section 3). We also make some remarks on them.
@@
    for some Lagrange multiplier $\lambda_{i}<0$ depending on the component. Here $\cdot$$_{+}$ is the nonnegative (positive) part function defined as $\lambda]_{+}:=\max\{\lambda,0\}$;
@@
    the equations (EL) continue to hold on a $\delta$-neighbourhood of the $\Omega_{i}$. Here the $\delta$-neighbourhood of $\Omega_{i}$ is defined as $\bigcup_{y\in\Omega_{i}}\left\{x\in\mathbb{R}^{3}\mid|x-y|<\delta\right\}$;
@@
    if $P(\rho)$ satisfies assumption (F4’), then $\rho$ satisfies the reduced Euler-Poisson equations (EP’);
@@
Due to the discussion in subsection 2.2, if $\widetilde{\rho}$ is a local minimizer over ${R}\left(\mathbb{R}^{3}\right)$ of $E_{J}(\rho)$ (or by Corollary 4.6, even just over ${R}_{0}(\mathbb{R}^{3})$), then we may define:
@@
where $\omega=\frac{J}{I\left(\rho\right)}$, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating binary star system (or star-planet system) such that $E(\rho,v)$ and $E_{J}(\rho)$ are local minima. Moreover, $({\rho}(t,x),v(t,x))$ solves (EP) with $V(t,x)=V_{\widetilde{\rho}}\left(R_{-\omega t}x\right)$.
@@
The “Lagrange multiplier” $\lambda_{i}$ in Parts (iv) and (v) here is also referred to as the “chemical potential” in McCann’s paper [31]. In statistical mechanics, the chemical potential corresponds to $\frac{\partial E}{\partial N}$, where $E$ is the (free) energy and $N$ is the particle number. See also [7].
@@
We retain the term $\bar{x}(\rho)$ here to be consistent with the form of the Euler-Lagrange equations in McCann’s paper [31, Section 6]. One may also extend Theorem 2.22 to cases with nonzero center of mass, for instance, via the translation mentioned in Remark 2.21.
@@
If we further assume the fluid particles move with bounded velocities, then $Y_{t}(y)$ will be a Lipschitz function of $t$ uniformly in $y$, and it is evident that (26) will be controlled by a multiple of |st||s-t|. Thus $\rho_{t}$ is a continuous function with respect to time $t$.
@@
On the other hand, if the topology on ${V}\left(\mathbb{R}^{3}\right)$ is such that the map taking $\boldsymbol{w}\in\mathbb{R}^{3}$ to $v(x):=\boldsymbol{w}\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous, then a converse to the statement above is also true: any local minimizer subject to the vector constraint $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$ also minimizes locally among the larger class of competitors with prescribed $J_{z}(\rho,v)=J$ [31, Remark 2.3]. This condition that $w\in\mathbb{R}^{3}\mapsto v(x):=w\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous is used in the following arguments. Due to [31, Remark 3.3], we can have a similar result to Proposition 2.17: if $\boldsymbol{J}$ is prescribed, the velocity which minimizes the energy is given by $\boldsymbol{v}(\rho,\boldsymbol{J})(x)=\boldsymbol{w}(\rho,\boldsymbol{J})\times x$, where $\boldsymbol{w}(\rho,\boldsymbol{J})\in\mathbb{R}$ is the unique angular velocity compatible with the given density $\rho$ and angular momentum $\boldsymbol{J}$. Notice the axis $\boldsymbol{w}$ of rotation may not coincide with the $z$-axis. One can show when $\rho$ is sufficiently close to $\rho^{\prime}$ in the sense of Wasserstein $L^{\infty}$ distance, $w(\rho,\boldsymbol{J})$ is sufficiently close to $w^{\prime}(\rho^{\prime},\boldsymbol{J^{\prime}})$. This condition that $w\in\mathbb{R}^{3}\mapsto v(x):=w\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous ensures that $\boldsymbol{v}(\rho,\boldsymbol{J})$ is sufficiently close to $\boldsymbol{v^{\prime}}(\rho^{\prime},\boldsymbol{J^{\prime}})$ in $V(\mathbb{R}^{3})$. In another word, local perturbations in ${R}_{0}\left(\mathbb{R}^{3}\right)$ can create local perturbations in ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$, which can help us to prove that any local minimizer subject to the vector constraint $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$ also minimizes locally among the larger class of competitors with prescribed $J_{z}(\rho,v)=J$, using similar arguments to [31, Remark 2.3 and Section 3].
@@
Now we can obtain a solution to (EP’) from the Euler-Lagrange equation (EL) thanks to Theorem 2.22 (iv). Conversely, for $\rho\in R_{0}(\mathbb{R}^{3})$ with N connected components to be a solution of (EP’), similar arguments as in subsection 2.2 shows that (EL) in Theorem 2.22 (iv) is necessary.
By the arguments of Auchmuty and Beals [5] or Li [26] or McCann [31], we know a constrained energy minimizer on $W_{m,J}$ exists and its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. Then such constrained energy minimizer is indeed a local minimizer due to Lemma 4.5 (ii), and hence a solution to (EP’) after translation.
@@
What’s more, after another translation, the center of mass of $\widetilde{\rho}$ can be 0 and is a local minimizer of $E_{J}(\rho)$. Let $v(x):=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\widetilde{\rho})}$, then the pair $(\widetilde{\rho},v)$ minimizes $E(\rho,v)$ locally on ${R}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ (thus on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right))$ subject to the constraint $J_{z}(\rho,v)=J$ or $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$. $\widetilde{\rho}$ satisfies reduced Euler-Poisson equations (EP’). Moreover, the uniformly rotating fluid $({\rho}(t,x),v(t,x))$ solves (EP) with $V(t,x)=V_{\widetilde{\rho}}\left(R_{-\omega t}x\right)$, here $({\rho}(t,x),v(t,x))=\left(\widetilde{\rho}\left(R_{-\omega t}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
@@
We point out that in McCann’s proof of Theorem 2.22, the discussion on how to transition from (EL) to (EP’) is not particularly detailed. One can see intuitively from subsection 2.2 that a refined discussion on differentiability is required. Therefore, in this section, we will go through the proof of Theorem 2.22, give new ideas for establishing differentiability (Part (vi) and Part (vii) of Theorem 2.22), and explain the differences between our proof and McCann’s.
@@
Given $s>0,h\in\mathbb{R}$ with $0<|h|\leq s$, let $t=\phi(s)$ and $t+l=\phi(s+h)$, then $h\rightarrow 0$ is equivalent to $l\rightarrow 0$ due to the continuity and monotonicity of $\phi$, and we also have $A^{\prime}(t)=s,A^{\prime}(t+l)=s+h$ since $\phi=\left(A^{\prime}\right)^{-1}$. Recall by (4) we have $A^{\prime}(s)s-A(s)=P(s)$. Therefore, we have
@@
Let $B(l)=lA^{\prime}(t)-A(t+l)$, then $B$ is differentiable and $B^{\prime}(l)=A^{\prime}(t)-A^{\prime}(t+l)\neq$ 0 if $l\neq 0$, since $A^{\prime}$ is strictly increasing. Notice $lA^{\prime}(t)+A(t)-A(t+l)=B(l)-B(0)$, by Cauchy’s mean value theorem [35, Theorem 5.9], $\exists\theta\in(0,1)$, such that
$(B(l)-B(0))A^{\prime\prime}(t+\theta l)=\left(A^{\prime}(t+l)-A^{\prime}(t)\right)B^{\prime}(\theta l)$
Since $l\neq 0$, $\left(A^{\prime}(t+l)-A^{\prime}(t)\right)\neq 0$, $B^{\prime}(\theta l)\neq 0$, then both (B(l)B(0))(B(l)-B(0)) and $A^{\prime\prime}(t+\theta l)$ are also not 0. Then
@@
By Lagrange’s mean value theorem [35, Theorem 5.8], $\exists\widetilde{\theta}\in(0,1)$, such that $A^{\prime}(t)-A^{\prime}(t+\theta l)=(-\theta l)A^{\prime\prime}(t+\widetilde{\theta}\theta l)$. Notice $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$ when $s>0$. Thus
@@
By Taylor’s theorem [35, Theorem 5.15], together with (F4’), we know $\exists\theta_{1}\in(0,1),\theta_{2}\in(0,1)$,
$P^{\prime}(t+\theta l)=P^{\prime}(t)+\theta lP^{\prime\prime}(t)+\frac{(\theta l)^{2}}{2}P^{(3)}(t)+\cdots+\frac{(\theta l)^{n}}{n!}P^{(n+1)}\left(t+\theta_{1}\theta l\right)=\frac{(\theta l)^{n}}{n!}P^{(n+1)}\left(t+\theta_{1}\theta l\right)\neq 0$
Similarly, we have
$P^{\prime}(t+\widetilde{\theta}\theta l)=\frac{(\widetilde{\theta}\theta l)^{n}}{n!}P^{(n+1)}\left(t+\theta_{2}\widetilde{\theta}\theta l\right)\neq 0$
Thus $\left|\frac{P^{\prime}(t+\widetilde{\theta}\theta l)}{P^{\prime}(t+\theta l)}\right|\leq 2$ when $|l|$ small enough. As $l\rightarrow 0$, we have $\frac{A^{\prime}(t)-A^{\prime}(t+\theta l)}{A^{\prime\prime}(t+\theta l)}\rightarrow 0$, and then $\frac{P\circ\phi(s+h)-P\circ\phi(s)}{h}\rightarrow t=\phi(s)$ as $h\rightarrow 0$. Thus, the derivative of $P\circ\phi$ is $\phi(s)$. By the same arguments we know $P\circ\phi$ is right differentiable at 0 with right derivatives $\phi(0)=$ 0. ∎
We skip the proofs of Part (i) and Part (viii) of Theorem 2.22; one can refer to McCann’s paper [31]. Part (ii) and Part (iii) are closely related to the context, so we revisit McCann’s proof here. The other parts of the proof are relatively new, either addressing details not mentioned by McCann (Parts (iv-v)) or providing new approaches to proving differentiability and explaining the necessity of modifying the original statements of the conclusions (Parts (vi-vii)).
@@
Due to Part (ii), it suffices to consider $\rho$ that locally minimizes $E_{J}(\rho)$ on $R(\mathbb{R}^{3})$. We will give proofs of “weak” versions of Parts (iv-v) (Proposition 5.7, Proposition 5.10 and Remark 5.11) later in subsection 5.2, that is, (EL) holds for almost all points in the region we consider with negative $\lambda_{i}$. Based on this “weak” version of (v), we obtain the global continuity in Part (iii) as the following: first notice that almost everywhere in the region $\{\rho>0\}$, (EL) holds true and $A^{\prime}(\rho)>0$ implies $\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}>0$. Therefore, on $\{\rho>0\}$, since $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ has compact support, we have
$A^{\prime}(\rho(x))=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}\leq V_{\rho}(x)+C$
This result, together with Hardy-Littlewood-Sobolev Inequality [8, Theorem 1.7] or Proposition B.1, can help to show the bound of $\|\rho\|_{L^{p}}$ and $\left\|V_{\rho}\right\|_{L^{\widetilde{p}}}$ with $p$ and $\widetilde{p}$ increasing alternately and reaching infinity in finite steps (bootstrap method, similar as [5, Lemma 3 and Theorem A]). Then $V_{\rho}$ is continuously differentiable thanks to Proposition B.3. Then on a $\delta$-neighbourhood of $\{\rho>0\}$, denoted by $\Omega_{\delta}$, which contains $\partial\{\rho>0\}$, we have $\rho=\phi\circ\left[W_{\rho}\right]_{+}$ a.e., where $W_{\rho}(x)=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}$ and $\phi=\left(A^{\prime}\right)^{-1}$ exists and continuous by Remark 2.6. It means $\rho$ has a continuous representative there. Notice the points in $\overline{\mathbb{R}^{3}\backslash\Omega_{\delta}}$ are the interior points of $\{\rho=0\}$, hence $\rho$ is also continuous in $\overline{\mathbb{R}^{3}\backslash\Omega_{\delta}}$. Thus, we conclude Part (iii).
@@
For Part (vi), notice in the region $\{\rho>0\}$, we have $\rho=\phi\circ W_{\rho}$ by Remark 2.23. Thanks to Part (iii) and $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\rho\in C_{c}^{0}\left(\mathbb{R}^{3}\right)$, where $C_{c}^{0}\left(\mathbb{R}^{3}\right)$ denotes the family of continuous functions with compact support. In particular, $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{q}\left(\mathbb{R}^{3}\right)$ for any $q>3$. By Proposition B.3, we know $V_{\rho}$ is continuous differentiable in $\mathbb{R}^{3}$. Since $P(\rho)$ satisfies assumption (F4), by Remark 2.7 we know $\phi\in C^{1}((0,\infty))$. Apply chain rule (one can refer to [38, Theorem 10.1.15]), we obtain $\rho\in C^{1}(\{\rho>0\})$. ∎